# Accuracy and stability of the accelerated multi-direct-forcing immersed boundary method

### v2

@@ Acknowledgment
TK received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (803553). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission. KS thanks Mr. S. Tsujimura for providing the data for the ice slurry flow in Figs. 14 and 15.
@@ A.1 Full algorithm for moving boundary problems
    If the error from the no-slip boundary condition is sufficiently small, then go to Step 5.
@@ Appendix B Volume force with iteration
$\tilde{\bm{F}}^{n}_{\ell}\simeq-\sum_{j=1}^{N}\lambda_{\rm max}{\bm{g}}^{n}_{\ell}(\bm{X}_{j})\Delta{V}_{j}=-\lambda_{\max}\bm{F}^{n}_{\ell}.$ (93)

### v3

@@ Acknowledgment
KS was supported by JSPS KAKENHI grant number JP23H01341. TK received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (803553). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission. KS thanks Mr. S. Tsujimura for providing the data for the ice slurry flow in Figs. 15 and 16.
@@ A.1 Full algorithm for moving boundary problems
    If the error of the no-slip boundary condition is sufficiently small, then go to Step 5.
@@ Appendix B Volume force with iteration
$\tilde{\bm{F}}^{n}_{\ell}\simeq-\sum_{j=1}^{N}\lambda_{\rm max}{\bm{g}}^{n}_{\ell}(\bm{X}_{j})\Delta{V}_{j}=\lambda_{\max}\bm{F}^{n}_{\ell}.$ (94)